OpenStudy (anonymous):

Which ordered pairs in the form (x, y) are solutions to the equation 7x – 5y = 28? Choose all answers that are correct. A. (−6, −14) B. (−1, −7) C. (4, 10) D. (7, 9)

OpenStudy (anonymous):

\(\Large\color{hotpink}{you~going~to~be~nice~and~let~me~help~?~or~naw}\)

OpenStudy (anonymous):

btw to make you feel better about everyone calling you pretty and such , you're not , you're rather unattractive and pretty hideous i might add , i hope that makes you day :)

OpenStudy (anonymous):

Look @DUDE..IM..A..DUCK you cant tell me what to do? Ok. i wasn't rude i was mad because im doing work and all i see is messages saying oh your hot or sexy or etc. And it gets annoying. Espically when you dont even know the person. And it gets annoying i rather be called ugly then be called sexy by strangers i dont know

OpenStudy (anonymous):

*Especially

OpenStudy (anonymous):

well then you're welcome

OpenStudy (anonymous):

Whatever. Ok? i dont got time for arguing i have important things in life.

OpenStudy (anonymous):

boys will be boys , don't pay any attention to em , if you dont like comments from people you can set your profile to private messaging

OpenStudy (anonymous):

i'm not trying to argue btw

OpenStudy (anonymous):

Im a new person on here. I never heard about this website my friend showed me it i decided well why not it will help me with school work. Look im asking you polite If you do not want to answer this question you may leave and go some other question. But if you do like to answer it your more then welcome to answer it.

OpenStudy (anonymous):

a and b are both right

OpenStudy (anonymous):

You know there negative?

OpenStudy (anonymous):

Yeah, but look at this. 1 x 1 = 1. -1 x 1 = -1 1 x -1 = -1 -1 x -1 = 1

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

id appreciate a medal , so thank you

OpenStudy (anonymous):

i appreciate the medal *

OpenStudy (anonymous):

Yep